OpenStudy (anonymous):

Looking at the first question, can you explain to me how you do this: About Emission Spectrums

OpenStudy (anonymous):

That is the PDF file

OpenStudy (anonymous):

E= h * f f= c/lambda so lambda = c/f and type of radiation can be determined from the graphics above the table